# Categorification of generic Su-Zhang Character Formula

### v1

@@ 2.5 $\mathfrak{g}_{-1}$-generic weights
To each regular dominant weight $\lambda\in\Lambda$ (or $L^{()}(\lambda)$) we associate, following [stroppel2012highest], its weight diagram as following.
@@
    If $\lambda$ is dominant, then $\lambda$ is totally disconnected.
  3. 3.
  4. 4.
    $\operatorname{Res}^{\mathfrak{g}}_{\mathfrak{g}_{\overline{0}}}K_{\geq 0}\bigl(L_{\overline{0}}(\lambda)\bigr)\;\cong\;\operatorname{Res}^{\mathfrak{g}}_{\mathfrak{g}_{\overline{0}}}K_{\leq 0}(L_{\bar{0}}(\lambda-2\rho_{\bar{1}}))\;\cong\;\bigoplus_{S\subset\Delta_{\overline{1}}^{+}}L_{\overline{0}}\Bigl(\lambda-\sum_{\alpha\in S}\alpha\Bigr).$
  5. 5.
3 Narrow Verma modules $N^{()}(\lambda)$
@@ 4 BGG resolutions
Let $\mathfrak{g}=\mathfrak{gl}(m)\oplus\mathfrak{gl}(n)$ and assume that $\lambda$ is regular dominant integral. Then there exists an exact sequence $BGG_{\bar{0}}(\lambda)$ of $\mathfrak{g}$-modules
$0\to M_{\overline{0}}(w_{0}\cdot\lambda)\to\bigoplus_{\begin{subarray}{c}w\in W\\ \ell(w)=\ell(w_{0})-1\end{subarray}}M_{\overline{0}}(w\cdot\lambda)\to\cdots\to\bigoplus_{\begin{subarray}{c}w\in W\\ \ell(w)=1\end{subarray}}M_{\overline{0}}(w\cdot\lambda)\to M_{\overline{0}}(\lambda)\to L_{\overline{0}}(\lambda)\to 0.$
@@
By Section 3, we obtain the following complex, denoted by $\mathrm{BGG}^{()}(\lambda)$:
@@
By Section 3, we obtain the hom of complex $K_{\geq 0}\mathrm{BGG}_{\bar{0}}(\lambda)\to\mathrm{BGG}^{()}(\lambda)\to K_{\leq 0}\mathrm{BGG}_{\overline{0}}(\lambda-2\rho_{\overline{1}})$. In particular, we obtain following commutative diagram.
@@
Theorem 4.3 ([su2007character, chmutov2014weyl]).
@@
$\operatorname{ch}L^{()}(\lambda)$ $=\frac{\displaystyle\prod_{\beta\in\Delta_{\overline{1}}^{()+}}(1+e^{-\beta})}{\displaystyle\prod_{\gamma\in\Delta_{\overline{0}}^{+}}(1-e^{-\gamma})}\sum_{w\in W}(-1)^{\ell(w)}\frac{e^{w\cdot\lambda}}{\displaystyle\prod_{\beta\in\Gamma_{\lambda}}(1+e^{-\beta})}$
$=\sum_{w\in W}(-1)^{\ell(w)}e^{w\cdot\lambda}\,\frac{\displaystyle\prod_{\beta\in\Delta_{\overline{1}}^{()+}\setminus\Gamma_{\lambda}}(1+e^{-\beta})}{\displaystyle\prod_{\gamma\in\Delta_{\overline{0}}^{+}}(1-e^{-\gamma})}$
$=\sum_{w\in W}(-1)^{\ell(w)}\frac{\operatorname{ch}M^{()}(w\cdot\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{\lambda}}(1+e^{-\beta})}.$
@@
$\operatorname{ch}N^{()}(\lambda)=e^{\lambda}\,\frac{\displaystyle\prod_{\beta\in\Delta_{\overline{1}}^{()+}\setminus\Gamma_{\lambda}}(1+e^{-\beta})}{\displaystyle\prod_{\gamma\in\Delta_{\overline{0}}^{+}}(1-e^{-\gamma})}\\ =\frac{\displaystyle\operatorname{ch}M^{()}(\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{\lambda}}(1+e^{-\beta})}.$
@@
$=\sum_{w\in W}(-1)^{\ell(w)}\frac{\operatorname{ch}M^{()}(w\cdot\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{\lambda}}(1+e^{-\beta})}$
$=\sum_{w\in W}(-1)^{\ell(w)}\sum_{S\subset\Delta_{\overline{1}}^{()+}\setminus\Gamma_{\lambda}}\operatorname{ch}M_{\overline{0}}\Bigl(w\cdot\lambda-\sum_{\alpha\in S}\alpha\Bigr).$
@@
Note that $w\Gamma_{\lambda}=\Gamma_{w\cdot\lambda}$ for all $w\in W$.
Since the characters $\operatorname{ch}M_{\overline{0}}\bigl(w\cdot\lambda-\sum_{\alpha\in S}\alpha\bigr)$ (for varying $w$) are linearly independent, which yields the desired identity and completes the proof. $\square$
@@
$\operatorname{ch}L^{()}(\lambda)=e^{\lambda}\,\frac{\displaystyle\prod_{\beta\in\Delta_{\overline{1}}^{()+}\setminus\Gamma_{\lambda}}(1+e^{-\beta})}{\displaystyle\prod_{\gamma\in\Delta_{\overline{0}}^{+}}(1-e^{-\gamma})}\\ =\frac{\displaystyle\operatorname{ch}M^{()}(\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{\lambda}}(1+e^{-\beta})}.$
Shunsuke Hirota
Department of Mathematics, Kyoto University
Kitashirakawa Oiwake-cho, Sakyo-ku, 606-8502, Kyoto
E-mail address: shun299509732@gmail.com

### v2

@@ 2.5 $\mathfrak{g}_{-1}$-generic weights
To each regular dominant weight $\lambda\in\Lambda$ (or $L^{()}(\lambda)$) we associate, following [6], its weight diagram as following.
@@
  3. 3.
    If $\lambda$ is dominant, then $\lambda$ is totally disconnected.
  4. 4.
Proof. (1) and (2) are immediate from the definition.
(3) If $\lambda$ is not totally disconnected, then there exist $i$ and $j$ such that $\lambda_{i}=\lambda_{m+j}=\lambda_{i+1}+1=\lambda_{m+j-1}+1$. Then $\lambda$ and $\lambda-(\varepsilon_{i}-\delta_{j})-(\varepsilon_{i}-\delta_{j-1})$ lie in different $\rho$–shifted Weyl chambers, contradicting $\mathfrak{g}_{-1}$–genericity.
(4) This follows from the fact that a Verma module has no nontrivial self-extension. $\square$
3 Narrow Verma modules $N^{()}(\lambda)$
Lemma 3.1 ([31]).
For $\lambda,\lambda^{\prime}\in\Lambda$, the following statements hold:
  1. 1.
    $\operatorname{ch}M^{()}(\lambda)=\operatorname{ch}M^{(n^{m})}(\lambda-2\rho_{\bar{1}})$;
@@ 3 Narrow Verma modules $N^{()}(\lambda)$
    $\dim\operatorname{Hom}(M^{()}(\lambda),M^{(n^{m})}(\lambda-2\rho_{\bar{1}}))=1$;
@@ 4 BGG resolutions
Definition 4.1.
Let $\mathcal{A}$ be an abelian category. We write $\operatorname{ch}\mathcal{A}$ for the category of chain complexes in $\mathcal{A}$, i.e. objects are sequences $C^{\bullet}=(\cdots\to C^{i-1}\xrightarrow{d^{i-1}}C^{i}\xrightarrow{d^{i}}C^{i+1}\to\cdots)$ in $\mathcal{A}$ with $d^{i}\circ d^{i-1}=0$, and morphisms are chain maps.
@@
Let $\mathfrak{g}=\mathfrak{gl}(m)\oplus\mathfrak{gl}(n)$ and let $\lambda$ be regular dominant integral. Then there exists an exact sequence $BGG_{\overline{0}}(\lambda)$ of $\mathfrak{g}$–modules of the form
$0\to M_{\overline{0}}(w_{0}\cdot\lambda)\to\bigoplus_{\ell(w)=\ell(w_{0})-1}M_{\overline{0}}(w\cdot\lambda)\to\cdots\to\bigoplus_{\ell(w)=1}M_{\overline{0}}(w\cdot\lambda)\to M_{\overline{0}}(\lambda)\to L_{\overline{0}}(\lambda)\to 0.$
Moreover, $BGG_{\overline{0}}(\lambda)$ is indecomposable.
@@
By Section 3, we obtain the hom of complex $K_{\geq 0}\mathrm{BGG}_{\bar{0}}(\lambda)\to K_{\leq 0}\mathrm{BGG}_{\overline{0}}(\lambda-2\rho_{\overline{1}})$.
We obtain the image denoted by $\mathrm{BGG}^{()}(\lambda)$:
In particular, we obtain following commutative diagram.
@@
Note that $w\Gamma_{\lambda}=\Gamma_{w\cdot\lambda}$ for all $w\in W$.
Theorem 4.4 ([36, 16]).
@@
$\operatorname{ch}L^{()}(\lambda)$ $=\frac{\displaystyle\prod_{\beta\in\Delta_{\overline{1}}^{()+}}(1+e^{-\beta})}{\displaystyle\prod_{\gamma\in\Delta_{\overline{0}}^{+}}(1-e^{-\gamma})}\sum_{w\in W}(-1)^{\ell(w)}\frac{e^{w\cdot\lambda}}{\displaystyle\prod_{\beta\in\Gamma_{w\cdot\lambda}}(1+e^{-\beta})}$
$=\sum_{w\in W}(-1)^{\ell(w)}e^{w\cdot\lambda}\,\frac{\displaystyle\prod_{\beta\in\Delta_{\overline{1}}^{()+}\setminus\Gamma_{w\cdot\lambda}}(1+e^{-\beta})}{\displaystyle\prod_{\gamma\in\Delta_{\overline{0}}^{+}}(1-e^{-\gamma})}$
$=\sum_{w\in W}(-1)^{\ell(w)}\frac{\operatorname{ch}M^{()}(w\cdot\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{w\cdot\lambda}}(1+e^{-\beta})}.$
@@
$\operatorname{ch}N^{()}(\lambda)=e^{\lambda}\,\frac{\displaystyle\prod_{\beta\in\Delta_{\overline{1}}^{()+}\setminus\Gamma_{w\cdot\lambda}}(1+e^{-\beta})}{\displaystyle\prod_{\gamma\in\Delta_{\overline{0}}^{+}}(1-e^{-\gamma})}\\ =\frac{\displaystyle\operatorname{ch}M^{()}(\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{w\cdot\lambda}}(1+e^{-\beta})}.$
@@
$=\sum_{w\in W}(-1)^{\ell(w)}\frac{\operatorname{ch}M^{()}(w\cdot\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{w\cdot\lambda}}(1+e^{-\beta})}$
$=\sum_{w\in W}(-1)^{\ell(w)}\sum_{S\subset\Delta_{\overline{1}}^{()+}\setminus\Gamma_{w\cdot\lambda}}\operatorname{ch}M_{\overline{0}}\Bigl(w\cdot\lambda-\sum_{\alpha\in S}\alpha\Bigr).$
@@
Since the subspaces $V_{w}:=\mathrm{span}\Bigl\{\operatorname{ch}M_{\overline{0}}\bigl(w\cdot\lambda-\sum_{\alpha\in S}\alpha\bigr)\ \Bigm|\ S\subset\Delta_{\overline{1}}^{()+}\setminus\Gamma_{w\cdot\lambda}\Bigr\}$ (for varying $w$) are linearly independent, this yields the desired identity and completes the proof. $\square$
@@
$\operatorname{ch}L^{()}(\lambda)=e^{\lambda}\,\frac{\displaystyle\prod_{\beta\in\Delta_{\overline{1}}^{()+}\setminus\Gamma_{w\cdot\lambda}}(1+e^{-\beta})}{\displaystyle\prod_{\gamma\in\Delta_{\overline{0}}^{+}}(1-e^{-\gamma})}\\ =\frac{\displaystyle\operatorname{ch}M^{()}(\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{w\cdot\lambda}}(1+e^{-\beta})}.$